# Distributed Hybrid Sketching for $\ell_{2}$-Embeddings

### v1

@@ IV-A Analysis of our approach
In the next proposition we further justify why we consider OSEs for our hybrid approach, other than the fact that they provide privacy benefits. The caveat though is that we require oversampling according to $\hat{\beta}$, which seems to be inevitable with sampling based OSEs. We also note that even though $\mathbf{\Pi}_{1}$ in Theorem 3 cannot be a non-OSE $\ell_{2}$-s.e., $\mathbf{\Pi}_{2}$ can be.
@@
Consider for instance the leverage score sketching matrix, which is not an OSE. In this case, the global properties that would need to be inferred is the orthonormal basis $\mathbf{U}$ of $\mathbf{A}$. For any fixed $\iota\in\mathbb{N}_{k}$, it is not possible to locally determine $\mathbf{U}_{\iota}$ (7) without any knowledge of $\mathbf{A}_{\{k\}}\backslash\{\mathbf{A}_{\iota}\}$, as the alignment of $\mathbf{U}_{\iota}$ with $\mathbf{I}_{N}$ depends on the entire data matrix $\mathbf{A}$. It is though possible to obtain such a subspace embedding by obtaining a (global) sampling scheme according to $\{\|\mathbf{A}_{(i)}\|_{2}^{2}\}_{i=1}^{N}$, though we would still require further rounds of communication. It is worthwhile investigating whether it is possible to devise a scheme which obtains an approximate global leverage score sampling sketch through local sampling, with few rounds of communication and restricted/limited information sharing.
IV-B Security in Distributed Local Sketching
Another major benefit of considering random Gaussian matrices is security they provide. In this work, security means that the nodes’ raw data $\mathbf{A}_{i}$ cannot be shared with nor recovered by the coordinator, or potential eavesdroppers. Under the assumption that $\mathbf{A}_{i}$ is randomly sampled from a distribution with finite variance, then the mutual information per symbol between $\mathbf{S}_{i}\mathbf{A}_{i}$ (sketch of $\mathbf{A}_{i}$ observed by the coordinator) and $\mathbf{A}_{i}$, has a logarithmic upper bound in terms of the variance; which approaches zero as $n$ increases or if $r$ is selected appropriately [62, 63, 64, 20]. This implies information-theoretic security and privacy of the local data blocks. Similar results were obtained in [19, 50] for $\mathbf{P}_{\{k\}}$ random unitary matrices, which made different assumptions on $\mathbf{A}_{i}$ and the distribution of $\mathbf{U}_{i}$.
Proposition 4.
IV-C Implications to Distributed Gradient Descent
An implication of [22, Thm. 3], is that as long as the $\ell_{2}$-s.e. property is satisfied by the iterative sketching [^2] By iterative sketching, we mean that at each iteration of the algorithm, we apply a new sketching matrix. (or sampling) technique which takes place, we converge in expectation to the optimal solution (1) at a rate inversely proportional to the number of iterations. To establish that our sketching technique (Figure 1) can be utilized to obtain the same convergence guarantee; it suffices to show that if every server performs a different sampling $\mathbf{\Omega}_{i}^{[t]}$ at each iteration $t$, i.e. a new global underlying sketching matrix $\widehat{\mathbf{S}}^{[t]}$ is applied at each iteration, the overall gradient is an unbiased estimate of the gradient $\nabla_{\mathbf{x}}L_{ls}(\mathbf{A},\mathbf{b};\mathbf{x}^{[t]})$, which we show this in Proposition 5.
Proposition 5.
Corollary 2 ([22, Thm. 3]).
Assume a bounded gradient on the least squares objective resulting from distributed GD with iterative sketching; according to Algorithm 1 with step-size $\eta_{t}=1/(\alpha t)$ for fixed $\alpha>0$. Then, after $T$ iterations:
It is worth mentioning that even though we directly invoke [22, Thm. 3] to give a convergence guarantee when the central server aggregates the gradients and performs distributed gradient descent, the formulation here is vastly different to that of previous works. In those works, the projections are performed on the aggregated data, which is the main contribution of this section, i.e. we showed that you do not need to aggregate the data to get a global embedding. Furthermore, this application to distributed GD can be viewed as an extension of the prior works, where now the data is meant to be kept private from the coordinator rather than the servers, similar to the FL convention. Additionally, here we require less overall storage, as the coordinator only needs to keep truck of the updated estimates $\tilde{\mathbf{g}}^{[t]}$ and ${\hat{\mathbf{x}}}^{[t+1]}={\hat{\mathbf{x}}}^{[t]}-\eta_{t}\cdot{\hat{\mathbf{x}}}^{[t]}$; for $\eta_{t}$ an appropriate step-size.
IV-D Experiments on Distributed Sketching
In the following experiments we considered the errors according to the $\ell_{2}$-s.e. error (5) and relative regression error (6), for $\mathbf{A}\in\mathbb{R}^{18000\times 40}$ following a $t$-distribution, similar to the experiment of Figure 2. We considered $R$ varying from 30% to 90% of $N$, for different values of $k$, and $\mathbf{P}_{i}$ rescaled Gaussian matrices. Through these experiments, we convey that the difference in error is small, while we save on computing on the transformed blocks $\{\mathbf{P}_{i}\mathbf{A}_{i}\}_{i=1}^{k}$ by a factor of $(k_{1}/k_{2})^{2}$ on the overall computation when we move from $k=k_{1}$ to a smaller $k=k_{2}$. By combining steps 1)1)1 ) and 2)2)2 ) of Algorithm 1, we require only $\mathcal{O}(rndk)$ operations by each server. Additionally, more accurate approximations were observed with higher $N$.
Figure 3: Normalized $\ell_{2}$-s.e. error (5), for varying $k$ and $R$.
Figure 4: Regression error (6), for varying $k$ and $R$.
@@ V Hybrid Sketching
In [20], the authors proposed sequentially applying two distinct sketches; where the first is analogous to Figure 1, to form what they call a “hybrid sketch”. An illustration od such a hybrid sketching procedure is portrayed in Figure 5. There are potential scenarios where such sketches are useful, e.g. there may be communication bottlenecks between the computational nodes and the central coordinator who wants to collect a description of the global dataset, though the coordinator wants to also compress it locally in order to save on computations which will then be done on the final sketch. Communication-efficiency is a key concern with all distributed algorithms, including federated ones [29]. As was also noted in [20], it might be computationally feasible for server nodes to sample as many data points as possible; e.g. $R/k$, and then reduce the dimension of the globally sampled data to the final sketch dimension ${\tilde{R}}$ using another sketch with better error properties.
By communicating larger local sketches (through $\widehat{\mathbf{S}}$, with a higher $R$), the coordinator obtains a better approximation (Theorem 2), and can then apply a sketch $\mathbf{\Phi}\in\mathbb{R}^{\rho R\times R}$ to get a more succinct representation of $\mathbf{A}$, for $\rho\in(0,1)$. The initial reduced dimension is $\mu N$ for $\mu\in(0,1)$, i.e. the global sketch dimension through $\widehat{\mathbf{S}}$ is $R=\lceil\mu N\rceil{}$. We denote the final target dimension of the global sketch by ${\tilde{R}}=\lceil\rho R\rceil{}$, i.e. ${\tilde{R}}=\lceil(\rho\mu)N\rceil{}$.
@@
[…]
@@
In our setting; by applying a global sketch $\mathbf{\Phi}$ with better global properties than $\widehat{\mathbf{S}}$, the coordinator will thus have a better sketch in hand, as it is now projecting the global dataset $\mathbf{A}$ into a random lower-dimensional subspace formed by all the local datasets. This is also more beneficial, computationally and in terms of communication and privacy, to simply delivering $\mathbf{A}_{\{k\}}$ to the coordinator who will apply a global sketch to the aggregated $\mathbf{A}$. The composite sketching matrix after applying
@@ V-A Embedding of Hybrid Sketching
Empirically, we observed that our approach is beneficial when $\mathbf{\Phi}$ is a Rademacher sketch, and that in the case when $\mathbf{\Phi}$ is an SRHT, we do not outperform mainstream sketching approaches. Furthermore, if the local $\mathbf{S}_{\{k\}}$ through $\widehat{\mathbf{S}}$ are SRHTs, the performance is very close to when $\mathbf{S}_{\{k\}}$ are Rademacher; and we have the benefit of accelerating the initial sketching procedure which is also done distributively. In the case when $N$ is relatively small, we also observed that selecting the local and the global sketching matrices to be a random Rademacher projection, leads to superior results. In conclusion, the most beneficial combination for large $N$ is when $\mathbf{S}_{\{k\}}$ are SRHTs, and $\mathbf{\Phi}$ is a Rademacher random matrix, which we also justify theoretically. All in all, the ideal choice for sketching matrices in our setting of Figure 5 and Algorithm 2, is to first apply SRHTs locally through $\mathbf{S}_{\{k\}}$, and then a Rademacher sketch on $\widehat{\mathbf{A}}$ through $\mathbf{\Phi}$.
@@
Fix the parameters $(\epsilon,\delta)\in(0,1)$ and $\mathbf{U}\in\mathbb{R}^{N\times d}$. Assume we are given sketching matrices $\mathbf{\Pi}_{1}\in\mathbb{R}^{R\times N}$ and $\mathbf{\Pi}_{2}\in\mathbb{R}^{{\tilde{R}}\times R}$, both satisfying (5) with parameters $\tilde{\delta}\coloneqq\delta/2$ and ${\tilde{\epsilon}}\coloneqq\sqrt{1+\epsilon}-1$ for $\mathsf{im}(\mathbf{U})$ and $\mathsf{im}(\mathbf{\Pi}_{1}\mathbf{U})$ respectively. Then, the hybrid sketching matrix $\hat{\mathbf{S}}\coloneqq\mathbf{\Pi}_{2}\mathbf{\Pi}_{1}$ satisfies (5) for $\mathsf{im}(\mathbf{U})$ with parameters $(\epsilon,\delta)$.
@@
Note that depending on the choices of $\mathbf{\Pi}_{1}$ and $\mathbf{\Pi}_{2}$, one can further optimize the parameters $({\tilde{\epsilon}},\tilde{\delta})$ to get superior results on either the sample or time complexity of the resulting sketch, though there will always be a trade-off between two. Furthermore, one can also extend the above result to consider the composition of more than just two sketching matrices.
@@
Experimentally (Sec. VI), we observed that the best performance in terms of the empirical $\ell_{2}$-s.e. error $\epsilon$ according to (5) occurs when the outer sketch is a random Rademacher matrix. Theoretically this choice is also sound, as it corroborates the facts conveyed in Table I. Specifically, Rademacher random projections require a smaller number of rows in comparison to the SRHT to guaranteed the same distortion. It is beneficial though in terms of complexity for the distributed servers to apply locally apply SRHTs. In summary, the outer sketching matrix $\mathbf{\Phi}$ has a greater affect on the overall error $\epsilon$. By reducing the complexity required for the application of $\mathbf{\Phi}$, we further accelerate the sketching of the global dataset $\mathbf{A}$, which was the main motivation for hybrid sketching [20]. In practice, this also resulted in better empirical distortion error compared to Gaussian sketches.
There are also practical implications to applying two distinct binary random matrices, as by what we propose; the entries of hybrid sketching matrix $\widetilde{\mathbf{S}}$ follow a shifted binomial distribution [^3] In the case where $\mathbf{S}_{\{k\}}$ and $\mathbf{\Phi}$ are Rademacher matrices. For $\mathbf{S}_{\{k\}}$ SRHTs, this is also true empirically. centered at 0. This is a consequence of the de Moivre–Laplace theorem, which can implies that $\widetilde{\mathbf{S}}$ can be viewed as a discrete approximation to a Gaussian random matrix, which are not easily realizable in practice. This gives further intuition and justification to our choices of $\mathbf{S}_{\{k\}}$ and $\mathbf{\Phi}$.
@@ V-B Local SRHTs with a Global Rademacher
Next, we focus on the ideal combination where $\mathbf{S}_{\{k\}}$ are SRHTs, and $\mathbf{\Phi}$ is a Rademacher sketching matrix. For this analysis, we make use of the sparse-SRHT [23]; a special case of our distributed sketching approach through $\widehat{\mathbf{S}}$. The reason we do so, is to avoid the dependence on the oversampling parameter $1/\beta$ in the sample complexity $R$ of Theorem 2, which considers a variety of random projection matrices; and not only the Randomized Hadamard Transform. It is also clear that unless we have a very low misestimation parameter $\beta$, then our $R$ is lower than that of [23]. Lastly, we note that in many algorithms; including the SRHT, to remove the dependence on $\beta$ an assumption on its value is made, e.g. $\beta=(2\ln(40Nd))^{-1}$ [11, 5]. Parameter $\beta$ is chosen such that an exact uniform sampling distribution results from the approximation of the flattened leverage scores, though it is significantly far from 1 for large $N$; which implies that the oversampling which takes place is relatively large.
Note that in the description of ‘Approach I’ in [23]; for practical reasons an additional permutation matrix is applied to $\mathbf{A}$, and it is assumed that the same indexed samples are selected from each locally projected dataset, i.e. $\mathbf{\Omega}_{i}$ is the same for all $i\in\mathbb{N}_{k}$. As was aforementioned, the global permutation does not affect the theoretical result, though it could have practical benefits. To alleviate the caveat that we allow each server to perform a distinct $\mathbf{\Omega}_{i}$, we can assume that a corresponding permutation is performed locally at each $\mathbf{A}_{i}$.
@@
For any $(\epsilon,\delta)\in(0,1)$, $k$ the number of blocks, and $\widehat{\mathbf{S}}\in\mathbb{R}^{R\times N}$ the sparse-SRHT, the $\ell_{2}$-s.e. property of (5) is satisfied with probability at least $1-\delta$ when
@@
The overall complexity of our application boils down to two matrix-matrix multiplications, by $\widehat{\mathbf{S}}$ by $\mathbf{\Phi}$, which are of complexity $\mathcal{O}(nd\log n)$ where $n=N/k$ for $k$ not too large, and $\mathcal{O}({\tilde{R}}Rd)=\mathcal{O}(N^{2}d\rho\mu^{2})$ respectively. The overall complexity of our approach is then $\mathcal{O}\big{(}Nd(\log(N/k)/k+N\rho\mu^{2})\big{)}$ which is asymptotically less than that of the SRHT when
@@
In terms of overall time complexity, not considering the fact that our computations are also carried out distributively, of our approach’s number of operations is $\mathcal{O}\big{(}Nd\big{(}\log(N/k)+N\mu^{2}\rho)\big{)}\big{)}$, while that of the Gaussian sketch is $\mathcal{O}(N{\tilde{R}}d)=\mathcal{O}(N^{2}d\rho\mu)$. By dropping the $\mathcal{O}$-notation, our approach is then superior when $\rho\mu(1-\mu)>\frac{\log(N/k)}{N}$ which holds true in general, unless one selects very small $\rho$ and $\mu$. In the case where $\rho$ and $\mu$ are very small though, it is likely that the required number of samples by the respective sketch (according to Table I) may not be satisfied. Furthermore, even for relative small compression factors of, e.g. $\rho=\mu=1/2$; our method is superior when $k>8$.

### v2

@@ IV-A Analysis of our approach
[…]
In the next proposition we further justify why we consider OSEs for our hybrid approach, other than the fact that they provide privacy benefits. The caveat though is that we require oversampling according to $\hat{\beta}$, which seems to be inevitable with sampling based OSEs. We also note that even though $\widehat{\mathbf{S}}$ in Theorem 3 cannot be a non-OSE $\ell_{2}$-s.e., $\mathbf{\Phi}$ can be.
@@
[…]
@@ V Hybrid Sketching
In [20], the authors proposed sequentially applying two distinct sketches, where the first is analogous to Fig. 1, to form what they call a “hybrid sketch”. An illustration of such a hybrid sketching procedure is portrayed in Fig. 3. There are potential scenarios where such sketches are useful, e.g. there may be communication bottlenecks between the computational nodes and the central coordinator who wants to collect a description of the global dataset, though the coordinator wants to also compress it locally in order to save on computations which will then be done on the final sketch. Communication-efficiency is a key concern with all distributed algorithms, including federated ones [29]. As was also noted in [20], it might be computationally feasible for server nodes to sample as many data points as possible; e.g. $R/k$, and then reduce the dimension of the globally sampled data to the final sketch dimension $\tilde{R}$} using another sketch with better error properties.
By communicating larger local sketches (through $\widehat{\mathbf{S}}$, with a higher $R$), the coordinator obtains a better approximation (Theorem 2), and can then apply a global sketch $\mathbf{\Phi}\in\mathbb{R}^{\rho R\times R}$ to get a more succinct representation of $\mathbf{A}$, for $\rho\in(0,1)$. The initial reduced dimension is $\mu N$ for $\mu\in(0,1)$, i.e. the global sketch dimension through $\widehat{\mathbf{S}}$ is $R=\lceil\mu N\rceil{}$. We denote the final target dimension of the global sketch by ${\tilde{R}}=\lceil\rho R\rceil{}$, i.e. ${\tilde{R}}=\lceil(\rho\mu)N\rceil{}$.
@@
$\mathsf{im}(\mathbf{A})\ \subseteq\ \mathbb{R}^{N}$$\mathsf{im}(\widehat{\mathbf{A}})\subseteq\ \mathbb{R}^{R}$${\mathbb{R}^{{\tilde{R}}}\ \supseteq\ \mathsf{im}(\widetilde{\mathbf{A}})}$$\scriptstyle{\widehat{\mathbf{S}}}$$\scriptstyle{\widetilde{\mathbf{S}}=\mathbf{\Phi}\circ\widehat{\mathbf{S}}}$$\scriptstyle{\mathbf{\Phi}}$
@@
In our setting, by applying a global sketch $\mathbf{\Phi}$ with better global properties than $\widehat{\mathbf{S}}$, the coordinator will thus have a better sketch in hand, as it is now projecting the global dataset $\mathbf{A}$ into a random lower-dimensional subspace formed by all the local datasets. This is also more beneficial, computationally and in terms of communication and privacy, to simply delivering $\mathbf{A}_{\{k\}}$ to the coordinator who will apply a global sketch to the aggregated $\mathbf{A}$. The composite sketching matrix after applying
@@ V-A Embedding of Hybrid Sketching
Empirically, we observed that our approach is beneficial when $\mathbf{\Phi}$ is a Rademacher sketch, and that in the case when $\mathbf{\Phi}$ is an SRHT, we do not outperform mainstream sketching approaches. Furthermore, if the local $\mathbf{S}_{\{k\}}$ through $\widehat{\mathbf{S}}$ are SRHTs, the performance is very close to when $\mathbf{S}_{\{k\}}$ are Rademacher, and we have the benefit of accelerating the initial sketching procedure which is also done distributively. In the case when $N$ is relatively small, we also observed that selecting the local and the global sketching matrices to be a random Rademacher projection, leads to superior results. In conclusion, the most beneficial combination for large $N$ is when $\mathbf{S}_{\{k\}}$ are SRHTs, and $\mathbf{\Phi}$ is a Rademacher random matrix, which we also justify theoretically. All in all, the ideal choice for sketching matrices in our setting of Fig. 3 and Algorithm 2, is to first apply SRHTs locally through $\mathbf{S}_{\{k\}}$, and then a Rademacher sketch on $\widehat{\mathbf{A}}$ through $\mathbf{\Phi}$.
@@
Fix the parameters $\epsilon,\delta\in(0,1)$ and $\mathbf{U}\in\mathbb{R}^{N\times d}$. Assume we are given sketching matrices $\mathbf{\Pi}_{1}\in\mathbb{R}^{R\times N}$ and $\mathbf{\Pi}_{2}\in\mathbb{R}^{{\tilde{R}}\times R}$, both satisfying (5) with parameters $\tilde{\delta}\coloneqq\delta/2$ and ${\tilde{\epsilon}}\coloneqq\sqrt{1+\epsilon}-1$ for $\mathsf{im}(\mathbf{U})$ and $\mathsf{im}(\mathbf{\Pi}_{1}\mathbf{U})$ respectively. Then, the hybrid sketching matrix $\hat{\mathbf{S}}\coloneqq\mathbf{\Pi}_{2}\mathbf{\Pi}_{1}$ satisfies (5) for $\mathsf{im}(\mathbf{U})$ with parameters $(\epsilon,\delta)$.
@@
[…]
@@
Experimentally (Sec. VII), we observed that the best performance in terms of the empirical $\ell_{2}$-s.e. error $\epsilon$ according to (5) occurs when the outer sketch is a random Rademacher matrix. Theoretically this choice is also sound, as it corroborates the facts conveyed in Table I. Specifically, Rademacher random projections require a smaller number of rows in comparison to the SRHT to guaranteed the same distortion. It is beneficial though in terms of complexity for the distributed servers to apply locally apply SRHTs. In summary, the outer sketching matrix $\mathbf{\Phi}$ has a greater affect on the overall error $\epsilon$. By reducing the complexity required for the application of $\mathbf{\Phi}$, we further accelerate the sketching of the global dataset $\mathbf{A}$, which was the main motivation for hybrid sketching [20]. In practice, this also resulted in better empirical distortion error compared to Gaussian sketches.
There are also practical implications to applying two distinct binary random matrices, as by what we propose, the entries of the hybrid sketching matrix $\widetilde{\mathbf{S}}$ follow a shifted binomial distribution [^1] In the case where $\mathbf{S}_{\{k\}}$ and $\mathbf{\Phi}$ are Rademacher matrices. For $\mathbf{S}_{\{k\}}$ SRHTs, this is also true empirically. centered at 0. This is a consequence of the de Moivre–Laplace theorem, which can implies that $\widetilde{\mathbf{S}}$ can be viewed as a discrete approximation to a Gaussian random matrix, which are not easily realizable in practice. This gives further intuition and justification to our choices of $\mathbf{S}_{\{k\}}$ and $\mathbf{\Phi}$.
@@ V-B Local SRHTs with a Global Rademacher
Next, we focus on the ideal combination where $\mathbf{S}_{\{k\}}$ are SRHTs, and $\mathbf{\Phi}$ is a Rademacher sketching matrix. For this analysis we make use of the sparse-SRHT [23], a special case of our distributed sketching approach through $\widehat{\mathbf{S}}$. The reason we do so, is to avoid the dependence on the oversampling parameter $1/\beta$ in the sample complexity $R$ of Theorem 2, which considers a variety of random projection matrices and not only the Randomized Hadamard Transform. It is also clear that unless we have a very low misestimation parameter $\beta$, then our $R$ is lower than that of [23]. Lastly, we note that in many algorithms; including the SRHT, to remove the dependence on $\beta$ an assumption on its value is made, e.g. $\beta=(2\ln(40Nd))^{-1}$ [11, 5]. Parameter $\beta$ is chosen such that an exact uniform sampling distribution results from the approximation of the flattened leverage scores, though it is significantly far from 1 for large $N$, which implies that the oversampling which takes place is relatively large.
Note that in the description of ‘Approach I’ in [23] for practical reasons an additional permutation matrix is applied to $\mathbf{A}$, and it is assumed that the same indexed samples are selected from each locally projected dataset, i.e. $\mathbf{\Omega}_{i}$ is the same for all $i\in\mathbb{N}_{k}$. As was aforementioned, the global permutation does not affect the theoretical result, though it could have practical benefits. To alleviate the caveat that we allow each server to perform a distinct $\mathbf{\Omega}_{i}$, we can assume that a corresponding permutation is performed locally at each $\mathbf{A}_{i}$.
@@
For any $\epsilon,\delta\in(0,1)$, $k$ the number of blocks, and $\widehat{\mathbf{S}}\in\mathbb{R}^{R\times N}$ the sparse-SRHT, the $\ell_{2}$-s.e. property of (5) is satisfied with probability at least $1-\delta$ when
@@
The overall complexity of our application boils down to two matrix-matrix multiplications, by $\widehat{\mathbf{S}}$ and $\mathbf{\Phi}$, which are of complexity $\mathcal{O}(nd\log n)$ where $n=N/k$ for $k$ not too large, and $\mathcal{O}({\tilde{R}}Rd)=\mathcal{O}(N^{2}d\rho\mu^{2})$ respectively. This complexity is then $\mathcal{O}\big(Nd(\log(N/k)/k+N\rho\mu^{2})\big)$ which is asymptotically less than that of the SRHT when
@@
In terms of overall time complexity, not considering the fact that our computations are also carried out distributively, of our approach’s number of operations is $\mathcal{O}\big(Nd\big(\log(N/k)+N\mu^{2}\rho)\big)\big)$, while that of the Gaussian sketch is $\mathcal{O}(N{\tilde{R}}d)=\mathcal{O}(N^{2}d\rho\mu)$. By dropping the $\mathcal{O}$-notation, our approach is then superior when $\rho\mu(1-\mu)>\frac{\log(N/k)}{N}$ which holds true in general, unless one selects very small $\rho$ and $\mu$. In the case where $\rho$ and $\mu$ are very small though, it is likely that the required number of samples by the respective sketch (according to Table I) may not be satisfied. Furthermore, even for relative small compression factors of, e.g. $\rho=\mu=1/2$, our method is superior when $k>8$.